OpenStudy (anonymous):

how to simplify underroot solution and how can we get the roundoff answer of any rootSquare.

OpenStudy (amistre64):

hmmm, maybe assume\[\sqrt{p}=\sqrt{n^2+\frac 1x}\] \[p=n^2+\frac 1x \]and solve for x

OpenStudy (amistre64):

i think i remembered that a little off \[\sqrt{p}=n+\frac{1}{x}\] \[p=n^2+\frac{2n}{x}+\frac{1}{x^2}\] ignore the 1/x^2 and solve for x

OpenStudy (amistre64):

spose we want say:\[\sqrt{123}\]that is close to 10^2 right? \[\sqrt{123}=10+\frac{1}{x}\] \[123=100+\frac{20}{x}\] \[\frac{20}{23}=x\]therefore\[\sqrt{123}=abt~10+\frac{23}{20}\]

OpenStudy (anonymous):

if i take 27

OpenStudy (amistre64):

lets try to define the x in general by this method: \[x=\frac{2n}{p-n^2}\]\[\sqrt{27}=5+\frac1x\]\[x=\frac{10}{27-25}=5\] \[\sqrt{27}~=abt~5.2\] we can keep refining i\[\sqrt{27}=5.2+\frac{1}{x}\] \[x=\frac{2(5.2)}{27-(5.2)^2}\]

OpenStudy (anonymous):

from where 5 comes

OpenStudy (amistre64):

5 is a guess. since 5^2 = 25 and 6^2 = 36, we know that sqrt(27) is somewhere between 5 and 6, so 5+1/x is a way to refine it

OpenStudy (anonymous):

this point i din't understand

OpenStudy (amistre64):

we could have well of started with sqrt(27) = 1 + 1/x and did some iterations x = 2/26; 1 + 26/2 = 14 let the new guess be 14 x = 28/(27-14^2) = -28/169 14 - 169/28 =abt 8 refine again

OpenStudy (amistre64):

if we can get a good guess to begin with, the refinements converge quicker

OpenStudy (amistre64):

 another method is to approximate the sqrt(x) function with a linear function by determing the slope at say x=25 y = sqrt(x) y' = 1/2sqrt(x) ; at x=25, y=1/10 using the point (25,5), and the slope of 1/10. we can determine an approximation for x=27 ... which is just 2 away from 25 y-5=1/10 (x-25) y-5=1/10 (27-25) y-5=1/10 (2) y = 5.2